OpenStudy (anonymous):

Simplify the expression –4(x – 7) + 3(x + 9). pls show me the work you did

OpenStudy (anonymous):

@Peach-es1234

OpenStudy (anonymous):

@countonme123

OpenStudy (anonymous):

@gibbs

OpenStudy (anonymous):

–4(x – 7) + 3(x + 9). =-4x +28+3x+18 =46-x

OpenStudy (anonymous):

@carsonce

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

:) anytime you can ask me .